OpenStudy (boldjon):

Aaron tracks the time it takes him to mow lawns by writing coordinate points relating x, the time in hours it takes to mow a lawn, and y, the size of the lawn in acres. Two of his points are (3, 1.5) and (5, 2.5). Which statement describes the slope of the line through these two points? A. It takes Aaron about 1 hour to mow 2 acres. B.Aaron’s rate for mowing lawns is 0.5 acres per hour. C.The ratio of acres to time is 5 acres to 3 hours. D.The rate to mow a lawn is 0.6 hours per acre.

OpenStudy (pawanyadav):

Slope is 2.5-1.5/(5-3)=1/2=0.5

OpenStudy (pawanyadav):

So which should be right opoption

OpenStudy (pawanyadav):

Can you tell..

OpenStudy (boldjon):

nope

OpenStudy (pawanyadav):

Slope = rate of change of y with x Here y =no. of acres And x=no. of hours And slope is 0.5 Can you guess now after watching the options carefully.

OpenStudy (boldjon):

yes....i can

OpenStudy (boldjon):

thanks a lot man